# Geometric Algebra Planes: Convex Implicit Neural Volumes

### v1

@@ A.7.1 Radiance Field Modeling
GA-Planes model uses feature dimensions $[d_{1},d_{2},d_{3}]=[36,24,8]$ and resolutions $[r_{1},r_{2},r_{3}]=[128,32,24]$. Multiresolution grids are not used for this task since density prediction can be achieved by a simpler architecture. The model size is 0.22 M. Tri-Planes model has the feature dimension $d_{2}=4$, and resolution $r_{2}=128$ resulting in a total number of parameters of 0.2 M. Note that we fix these sizes across (non/semi)convex formulations, which causes slight variations in the size of the decoder, however, the grids constitute the most number of parameters, making this effect negligible.

### v2

@@ A.7.1 Radiance Field Modeling
GA-Planes model uses feature dimensions $[d_{1},d_{2},d_{3}]=[36,24,8]$ (with $\odot$) or $[d_{1},d_{2},d_{3}]=[25,25,8]$ (with $\circ$) and resolutions $[r_{1},r_{2},r_{3}]=[128,32,24]$. Multiresolution grids are not used for this task since density prediction can be achieved by a simpler architecture. The model size is 0.22 M. Tri-Planes model has the feature dimension $d_{2}=4$, and resolution $r_{2}=128$ resulting in a total number of parameters of 0.2 M. Note that we fix these sizes across (non/semi)convex formulations, which causes slight variations in the size of the decoder, however, the grids constitute the most number of parameters, making this effect negligible.